OpenStudy (anonymous):

what is the solution to the equation -6z + 1 = 13 a. z = -6 b. z = -2 c. z = 2 d. z = 6

OpenStudy (sleepyhead314):

Hello @taylorsuzanne505 and Welcome to OpenStudy! :) to solve for z in this equation, we want to get z alone so first try subtracting 1 from both sides -6z + 1 = 13 -1 -1 ------------ what would you get? :)

OpenStudy (anonymous):

im confused already, sorry. what exact number am i subtracting from? 13 & 1?

OpenStudy (anonymous):

@sleepyhead314

OpenStudy (sleepyhead314):

yes, 13 and 1 1 - 1 = 0 and 13 - 1 = 12 so you can get -6z + 1 = 13 -1 -1 ------------ -6z = 12 do you see that? :)

OpenStudy (anonymous):

okay, so. z would be -6?

OpenStudy (sleepyhead314):

not quite :) from -6z = 12 we should divide both sides by -6 /-6 /-6 so z = 12 divided by -6 = ?

OpenStudy (anonymous):

Z = -2

OpenStudy (sleepyhead314):

that's what I got :)

OpenStudy (anonymous):

can you help with a few more? these kind of problems confuse me.

OpenStudy (sleepyhead314):

of course :)

OpenStudy (anonymous):

okay one sec. let me post it.

OpenStudy (anonymous):

If 4 + 8a = 4, what is the value of a. a. -8 b. -1 c. 0 d. 1

OpenStudy (sleepyhead314):

try subtracting 4 from both sides first :)

OpenStudy (anonymous):

4 - 4 = 0 8 - 4 = 4

OpenStudy (sleepyhead314):

not quite, it's more like 4 + 8a = 4 -4 -4 ----------- 8a = 0

OpenStudy (anonymous):

i see now

OpenStudy (anonymous):

so its 0?

OpenStudy (sleepyhead314):

yep :)

OpenStudy (sleepyhead314):

sorry but I gtg now you can close this question and ask a new one if you have any more questions and someone else will be able to come and help you :)